# Brauer groups of abelian varieties over fields of finite characteristic

### v1

@@ Introduction
see [P24, Theorem A.1]. In contrast with the $\ell$-adic case, the group ${\rm H}^{3}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}]$ can be infinite. A phenomenon specific to finite characteristic is the existence of a natural commutative quasi-algebraic group over $k$ whose group of $k$-points is ${\rm H}^{3}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}]$. This goes back to Milne [Mil76, Cor. 2.7 (a)], see [Ber81, Cor. 2.10], [IR83, Lemme IV.3.2.2], [Mil86, Lemma 1.8]. Quasi-algebraic groups were introduced by Serre [Ser60, §1.2] as equivalence classes of algebraic groups with a common set of $k$-points, where two groups $G$ and $H$ are called equivalent if there is a morphism $G\to H^{(p^{n})}$ which is a bijection on $k$-points.
When $X=A$ is an abelian variety, the attached quasi-algebraic group is connected, see [YY, Thm. 1.5] based on [IR83, Thm. II.3.4] (“survie du cœur”). Hence this is a commutative connected unipotent quasi-algebraic group, which we denote by ${\bf U}_{A}$. In particular, the group of $k$-points ${\bf U}_{A}(k)\cong{\rm H}^{3}(A,{\mathbb{Z}}_{p}(1))[p^{\infty}]$, the dimension ${\rm{dim}}({\bf U}_{A})$, and the $p$-exponent $\operatorname{\mathnormal{p}-exp}({\bf U}_{A})$ (the least integer $m$ such that $p^{m}{\bf U}_{A}=0$) are all well-defined. By general theory [Ser59, VII.10], ${\bf U}_{A}$ is isogenous to a product of truncated Witt (quasi-algebraic) groups $\prod_{i=1}^{r}{\bf W}_{n_{i}}$, for some positive integers $n_{1}\leqslant\ldots\leqslant n_{r}$. We have
@@
where $g={\rm{dim}}(A)$ and $m_{\lambda}$ is the multiplicity of slope $\lambda$ in ${\rm H}^{2}_{\rm cris}(A/W)$. In particular, ${\rm{dim}}({\bf U}_{A})$ depends only on the slopes of $A$ and thus is an invariant of the isogeny class of $A$. The meaning of the terms of (2) is the following: ${\rm{dim}}_{k}{\rm H}^{0}(A,\Omega^{2}_{A/k})=g(g-1)/2$ is the dimension of the tangent space of the formal Brauer group $\widehat{{\rm{Br}}}(A)$, the sum is the dimension of the maximal quotient of $\widehat{{\rm{Br}}}(A)$ of finite height, and ${\rm{dim}}({\bf U}_{A})$ is the dimension of the unipotent part of $\widehat{{\rm{Br}}}(A)$. Thus (2) implies the smoothness of the formal Brauer group of any abelian variety over $k$. Note that smoothness of $\widehat{{\rm{Br}}}(A)$ is a particular case of a general criterion of smoothness of Artin–Mazur formal groups obtained by the first named author in [Gr25b].
@@
The work on this paper started when A.S. was a Shapiro Visitor in Pennsylvania State University. He is very grateful to Yuri Zarhin for many helpful discussions without which this paper would have never been written. L.G. received funding from the EU Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement 101126554. The work of Y.Y. was supported by the London School of Geometry and Number Theory through the Engineering and Physical Sciences Research Council grant EP/S021590/1. The authors are grateful to the Bernoulli Center at EPFL, where a large part of this work was carried out, for excellent working conditions and support. We would like to thank James Borger, George Boxer, Daniel Bragg, Vladimir Drinfeld, Mikhail Kapranov, and Alexander Petrov for their questions and helpful discussions.
@@ 3.1 The dimension of ${\bf U}_{A}$
Proof. It is enough to check that the dimension of the formal Brauer group $\widehat{{\rm{Br}}}(A)$ equals the dimension of the tangent space to $\widehat{{\rm{Br}}}(A)$ at 0, which is ${\rm H}^{2}(A,{\cal O}_{A})$, see [AM77, Cor. II.2.4]. As we explain in Section 3.2 below, this equality is equivalent to formula (15). $\Box$
3.2 ${\bf U}_{A}$ and the formal Brauer group of AA
We now give a different computation of ${\rm{dim}}({\bf U}_{A})$, which exploits the relation between ${\bf U}_{A}$ and the formal Brauer group of $A$. For a smooth proper variety $X$ over $k$, the formal Brauer group $\widehat{\mathrm{Br}}(X)$ is the functor parameterising deformations of the trivial class in ${\rm{Br}}(X)$. For an Artinian $k$-algebra $R$ one defines
@@ 3.2 ${\bf U}_{A}$ and the formal Brauer group of AA
$(1)$ If ${\rm{Pic}}_{X/k}$ is smooth, then $\widehat{\mathrm{Br}}(X)$ is representable by a connected formal group of finite type with tangent space ${\rm H}^{2}(X,\mathcal{O}_{X})$.
@@
$(3)$ If $\widehat{{\rm{Br}}}(X)$ is representable, then its covariant Dieudonné module is isomorphic to ${\rm H}^{2}(X,W\mathcal{O}_{X})$.
Proof. Parts (1) and (3) are proved in [AM77], see Proposition II.1.8, Corollaries II.2.4 and II.4.3. Part (2) is a consequence of work of Ekedahl [Eke85]. For a different proof, see [Gr25b, Corollary 3.17]. $\Box$
From now on we assume that $X$ is an abelian variety $\pi\colon A\to{\rm{Spec}}(k)$, so that the conclusions of Proposition 3.4 (1), (2), (3) are true. In particular, $\widehat{\mathrm{Br}}(A)$ is a smooth connected formal group of dimension $h^{0,2}(A)=g(g-1)/2$. Recall [Ha, (28.3.2)] that a smooth connected formal group of finite type is an extension of a smooth connected formal group of finite height by a smooth unipotent formal group, in a unique way. Let us denote by $\widehat{U}_{A}$ the unipotent part of $\widehat{\mathrm{Br}}(A)$, and by $\widehat{P}_{A}$ the maximal finite height quotient of $\widehat{\mathrm{Br}}(A)$. By construction, $\widehat{U}_{A}$ is the reduced formal subgroup of $\widehat{\mathrm{Br}}(A)[p^{n}]$ for large enough $n$.
By a theorem of Bragg and Olsson [BO21, Cor. 1.6] we know that $R^{2}\pi_{*}\mu_{p^{n},A}$ is represented by a group $k$-scheme. Let us denote by ${\mathcal{U}}_{A,n}$ its reduced connected component. The exact sequence
$1\to\mu_{p^{n},A}\to\mu_{p^{n+1},A}\to\mu_{p,A}\to 1$
@@
The formal group $\widehat{U}_{A}$ is canonically isomorphic to the formal completion of ${\mathcal{U}}_{A}$ at the identity.
Proof. A more general statement is proved in [BO21, version 1, Prop. 10.7]. We spell out the argument in our case. Recall that the formal completion functor maps a presheaf $\mathcal{F}$ over ${\rm{Spec}}(k)_{\rm fppf}$ to the functor
@@
from the category of Artinian $k$-algebras to the category of abelian groups. Since the projection $R\to R_{\textnormal{red}}$ has a canonical retraction, the formal completion functor is exact. Pushing forward the Kummer sequence along $\pi:A\to{\rm{Spec}}(k)$ we obtain the short exact sequence of fppf sheaves
$0\to R^{1}\pi_{*}\mathbb{G}_{m,A}/p^{n}\to R^{2}\pi_{*}\mu_{p^{n},A}\to R^{2}\pi_{*}\mathbb{G}_{m,A}[p^{n}]\to 0,$
where $R^{1}\pi_{*}\mathbb{G}_{m,A}/p^{n}$ is representable by an étale group scheme, by the theory of the Picard scheme [CTS21, p. 121]. Therefore, taking the formal completion we get an isomorphism
$(R^{2}\pi_{*}\mu_{p^{n},A})^{\wedge}\cong(R^{2}\pi_{*}\mathbb{G}_{m,A}[p^{n}])^{\wedge}.$ (18)
Notice that the right-hand group is by definition $\widehat{\mathrm{Br}}(A)[p^{n}]$. Take $n$ large enough so that $\widehat{U}_{A}$ is the reduced formal subgroup of $\widehat{\mathrm{Br}}(A)[p^{n}]$, and ${\mathcal{U}}_{A}={\mathcal{U}}_{A,n}$. Then (18) gives a desired isomorphism $({\mathcal{U}}_{A})^{\wedge}\cong\widehat{U}_{A}$. $\Box$
By the Kummer sequence, the quasi-algebraic group $({\mathcal{U}}_{A})^{\rm perf}$ associated to ${\mathcal{U}}_{A}$ is isogenous to ${\bf U}_{A}$. Thus from the above discussion, we deduce the equalities:
${\rm{dim}}({\bf U}_{A})={\rm{dim}}(\widehat{U}_{A})={\rm{dim}}(\widehat{\mathrm{Br}}(A))-{\rm{dim}}(\widehat{P}_{A})=\frac{g(g-1)}{2}-{\rm{dim}}(\widehat{P}_{A}).$
It remains to compute ${\rm{dim}}(\widehat{P}_{A})$. By Proposition 3.4 (3), the covariant Dieudonné module of $\widehat{P}_{A}$ is isogenous to ${\rm H}^{2}(A,W\mathcal{O}_{A})$. The latter is isogenous to ${\rm H}^{2}_{\rm cris}(A/W)_{[0,1)}$ by the theory of the slope spectral sequence. It is well-known (see, e.g. [AM77, Proof of Cor. III.3.4]) that
${\rm{dim}}(\widehat{P}_{A})=\sum_{0\leqslant\lambda<1}m_{\lambda}(1-\lambda),$
where $\lambda$ is a slope of ${\rm H}^{2}_{\rm cris}(A/W)$ of multiplicity $m_{\lambda}$, see Lemma 3.6 below. This gives another proof of Theorem 3.1.
@@
Let $G$ be a $p$-divisible group. Let $M(G)$ be the covariant Dieudonné module of $G$. Then we have

### v2

@@ Introduction
see [P24, Theorem A.1]. In contrast with the $\ell$-adic case, the group ${\rm H}^{3}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}]$ can be infinite. A phenomenon specific to finite characteristic is the existence of a natural commutative quasi-algebraic group over $k$ whose group of $k$-points is ${\rm H}^{3}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}]$. Its connected component, which is unipotent, will be denoted by ${\bf U}_{X}$. This goes back to Milne [Mil76, Cor. 2.7 (a)], see [Ber81, Cor. 2.10], [IR83, Lemme IV.3.2.2], [Mil86, Lemma 1.8]. Quasi-algebraic groups were introduced by Serre [Ser60, §1.2] as equivalence classes of algebraic groups with a common set of $k$-points, where two groups $G$ and $H$ are called equivalent if there is a morphism $G\to H^{(p^{n})}$ which is a bijection on $k$-points.
When $X=A$ is an abelian variety, the attached quasi-algebraic group is connected and so is equal to ${\bf U}_{A}$, see [YY, Thm. 1.5] based on [IR83, Thm. II.3.4] (“survie du cœur”). In particular, the group of $k$-points ${\bf U}_{A}(k)\cong{\rm H}^{3}(A,{\mathbb{Z}}_{p}(1))[p^{\infty}]$, the dimension ${\rm{dim}}({\bf U}_{A})$, and the $p$-exponent $\operatorname{\mathnormal{p}-exp}({\bf U}_{A})$ (the least integer $m$ such that $p^{m}{\bf U}_{A}=0$) are all well-defined. By general theory [Ser59, VII.10], ${\bf U}_{A}$ is isogenous to a product of truncated Witt (quasi-algebraic) groups $\prod_{i=1}^{r}{\bf W}_{n_{i}}$, for some positive integers $n_{1}\leqslant\ldots\leqslant n_{r}$. We have
@@
where $g={\rm{dim}}(A)$ and $m_{\lambda}$ is the multiplicity of slope $\lambda$ in ${\rm H}^{2}_{\rm cris}(A/W)$. In particular, ${\rm{dim}}({\bf U}_{A})$ depends only on the slopes of $A$ and thus is an invariant of the isogeny class of $A$. The meaning of the terms of (2) is the following: ${\rm{dim}}_{k}{\rm H}^{0}(A,\Omega^{2}_{A/k})=g(g-1)/2$ is the dimension of the tangent space of the formal Brauer group $\widehat{{\rm{Br}}}(A)$, the sum is the dimension of the maximal quotient of $\widehat{{\rm{Br}}}(A)$ of finite height, and ${\rm{dim}}({\bf U}_{A})$ is the dimension of the unipotent part of $\widehat{{\rm{Br}}}(A)$, see Section 3.2. Thus (2) implies the smoothness of the formal Brauer group of any abelian variety over $k$ (Corollary 3.3). Note that smoothness of $\widehat{{\rm{Br}}}(A)$ is a particular case of a general criterion of smoothness of Artin–Mazur formal groups obtained by the first named author in [Gr25b].
@@
The work on this paper started when A.S. was a Shapiro Visitor in Penn State. He is very grateful to Yuri Zarhin for many helpful discussions without which this paper would have never been written. L.G. received funding from the EU Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement 101126554. The work of Y.Y. was supported by the London School of Geometry and Number Theory through the Engineering and Physical Sciences Research Council grant EP/S021590/1. The authors are grateful to the Bernoulli Center at EPFL, where a large part of this work was carried out, for excellent working conditions and support. We would like to thank James Borger, George Boxer, Daniel Bragg, Marco D’Addezio, Vladimir Drinfeld, Mikhail Kapranov, and Alexander Petrov for their questions and helpful discussions.
@@ 3.1 The dimension of ${\bf U}_{A}$
Proof. By the dimension formula (15), this follows from Proposition 3.7 below. $\Box$
3.2 Relation to the formal Brauer group
Let $X$ be a smooth proper variety over $k$, and let ${\bf U}_{X}$ be the connected component of the quasi-algebraic group over $k$ whose group of $k$-points is ${\rm H}^{3}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}]$ (as discussed in the introduction).
In this section we relate ${\bf U}_{X}$ to the formal Brauer group of $X$. The formal Brauer group $\widehat{\mathrm{Br}}(X)$ is the functor parameterising deformations of the trivial class in ${\rm{Br}}(X)$. For an Artinian $k$-algebra $R$ one defines
@@ 3.2 Relation to the formal Brauer group
$(1)$ If the Picard scheme ${\bf Pic}_{X/k}$ is smooth, then $\widehat{\mathrm{Br}}(X)$ is representable by a connected formal group of finite type with tangent space ${\rm H}^{2}(X,\mathcal{O}_{X})$.
$(2)$ If $\widehat{{\rm{Br}}}(X)$ is representable, then its Cartier module of typical curves is isomorphic to ${\rm H}^{2}(X,W\mathcal{O}_{X})$.
@@
Proof. Parts (1) and (2) are proved in [AM77], see Proposition II.1.8, Corollaries II.2.4 and II.4.3. Part (3) is a consequence of work of Ekedahl [Eke85]. For a different proof, see [Gr25b, Cor. 3.17]. $\Box$
Suppose that ${\bf Pic}_{X/k}$ is smooth, so that $\widehat{\mathrm{Br}}(X)$ is representable, and let $\widehat{\mathrm{Br}}(X)_{\rm red}$ be the reduced subgroup of $\widehat{\mathrm{Br}}(X)$. Recall [Ha, (28.3.2)] that a smooth connected formal group of finite type is an extension of a smooth connected formal group of finite height by a smooth unipotent formal group, in a unique way. Let us denote by $\widehat{U}_{X}$ the unipotent part of $\widehat{\mathrm{Br}}(X)_{\rm red}$, and by $\widehat{P}_{X}$ the maximal finite height quotient of $\widehat{\mathrm{Br}}(X)_{\rm red}$. By construction, $\widehat{U}_{X}$ is the reduced formal subgroup of $\widehat{\mathrm{Br}}(X)[p^{n}]$ for large enough $n$.
Let $\pi\colon X\to{\rm{Spec}}(k)$ be the structure morphism. By a theorem of Bragg and Olsson [BO21, Cor. 1.6] we know that $R^{2}\pi_{*}\mu_{p^{n},X}$ is represented by a group $k$-scheme. Let us denote by ${\mathcal{U}}_{X,n}$ its reduced connected component. The exact sequence
$1\to\mu_{p^{n},X}\to\mu_{p^{n+1},X}\to\mu_{p,X}\to 1$
@@
The formal group $\widehat{U}_{X}$ is canonically isomorphic to the formal completion of ${\mathcal{U}}_{X}$ at the identity.
Proof. A more general statement is proved in [BO21, version 1, Prop. 10.7]. Recall that the formal completion functor maps a presheaf $\mathcal{F}$ over ${\rm{Spec}}(k)_{\rm fppf}$ to the functor
@@
from the category of Artinian $k$-algebras to the category of abelian groups. Since the projection $R\to R_{\textnormal{red}}$ has a canonical section, the formal completion functor is exact. Pushing forward the Kummer sequence along $\pi:X\to{\rm{Spec}}(k)$ we obtain the short exact sequence of fppf sheaves
$0\to R^{1}\pi_{*}\mathbb{G}_{m,X}/p^{n}\to R^{2}\pi_{*}\mu_{p^{n},X}\to R^{2}\pi_{*}\mathbb{G}_{m,X}[p^{n}]\to 0,$
where $R^{1}\pi_{*}\mathbb{G}_{m,X}/p^{n}$ is representable by an étale group scheme, by the theory of the Picard scheme [CTS21, p. 121]. Therefore, taking the formal completion we get an isomorphism
$(R^{2}\pi_{*}\mu_{p^{n},X})^{\wedge}\cong(R^{2}\pi_{*}\mathbb{G}_{m,X}[p^{n}])^{\wedge}.$ (18)
Notice that the right-hand group is by definition $\widehat{\mathrm{Br}}(X)[p^{n}]$. Take $n$ large enough so that $\widehat{U}_{X}$ is the reduced formal subgroup of $\widehat{\mathrm{Br}}(X)[p^{n}]$, and such that ${\mathcal{U}}_{X}={\mathcal{U}}_{X,n}$. Then (18) gives a desired isomorphism $({\mathcal{U}}_{X})^{\wedge}\cong\widehat{U}_{X}$. $\Box$
By the Kummer sequence, the quasi-algebraic group $({\mathcal{U}}_{X})^{\rm perf}$ attached to ${\mathcal{U}}_{X}$ is isogenous to ${\bf U}_{X}$. From Lemma 3.5 we obtain
${\rm{dim}}({\bf U}_{X})={\rm{dim}}({\mathcal{U}}_{X})={\rm{dim}}(\widehat{U}_{X})={\rm{dim}}(\widehat{\mathrm{Br}}(X)_{\rm red})-{\rm{dim}}(\widehat{P}_{X}).$ (19)
Let us compute ${\rm{dim}}(\widehat{P}_{X})$. By Proposition 3.4 (2), the Cartier module of typical curves of $\widehat{\mathrm{Br}}(X)_{\rm red}$ is isomorphic to ${\rm H}^{2}(X,W\mathcal{O}_{X})$. It follows that, after inverting $p$, the covariant Dieudonné module of $\widehat{P}_{X}$ is isomorphic to ${\rm H}^{2}(X,W\mathcal{O}_{X})[1/p]$, hence it is isogenous to ${\rm H}^{2}_{\rm cris}(X/W)_{[0,1)}$ by the theory of the slope spectral sequence. It is then well-known (see, e.g. [AM77, Proof of Cor. III.3.4]) that
${\rm{dim}}(\widehat{P}_{X})=\sum_{0\leqslant\lambda<1}m_{\lambda}(1-\lambda),$ (20)
where $\lambda$ is a slope of ${\rm H}^{2}_{\rm cris}(X/W)$ of multiplicity $m_{\lambda}$, see Lemma 3.6 below.
@@
Let $G$ be a $p$-divisible group. Let $M(G)$ be the covariant Dieudonné module of $G$. We have
@@
Proposition 3.7
Let $X$ be a proper smooth variety over $k$ such that the Picard scheme ${\bf Pic}_{X/k}$ is smooth. Then the formal Brauer group $\widehat{\mathrm{Br}}(X)$ is smooth if and only if
$h^{0,2}(X)={\rm{dim}}({\bf U}_{X})+\sum_{\lambda}m_{\lambda}(1-\lambda),$
where $m_{\lambda}$ is the multiplicity of the slope $\lambda$ in ${\rm H}^{2}_{\rm cris}(X/W)_{[0,1)}$.
Proof. By Proposition 3.4 (1) the tangent space of $\widehat{\mathrm{Br}}(X)$ has dimension $h^{0,2}(X)={\rm{dim}}({\rm H}^{2}(X,\mathcal{O}_{X}))$. Thus $\widehat{\mathrm{Br}}(X)$ is smooth if and only if ${\rm{dim}}(\widehat{\mathrm{Br}}(X)_{\rm red})=h^{0,2}(X)$. It remains to apply (19) and (20). $\Box$
This gives another proof of Theorem 3.1 assuming smoothness of the formal Brauer group of abelian varieties.
When $X$ is a surface such that ${\bf Pic}_{X/k}$ is smooth, then $\widehat{\mathrm{Br}}(X)$ is smooth [AM77, p. 190]. In this case Proposition 3.7 follows from [Mil75, Remark 7.4].